diaddar8873:

Alex wants a new computer. Which option is the most financially responsible decision for the first two years

diaddar8873:

Determine the costs of the four computer options. The total cost which is the smallest will be the financially responsible decision. Calculate for 2 years (24 months), since his plan is to replace the computer in two years. Rental - $75 a month; once yearly maintenance for $400.00 The rental is $75 a month for 24 months, plus two $400 maintenance fees. Total cost = ($75 × 24) + ($400 × 2) = $2,600 Rental - $120 a month; free yearly maintenance The rental is $120 a month for 24 months with no maintenance fees. Total cost = $120 × 24 = $2,880 Purchase - $2,000; $700 two-year maintenance agreement The purchase is $2,000 plus $700 maintenance fee. Total cost = $2,000 + $700 = $2,700 Purchase - $1,900; once yearly maintenance for $400.00 The purchase is $1,900 plus two $400 maintenance fees. Total cost = $1,900 + ($400 × 2) = $2,700 Therefore, the financially responsible decision is the rental - $75 a month; once yearly maintenance for $400.00.